OpenStudy (anonymous):

Let f(x) = 5x + 12. Find f–1(x) 12x + 5 1/5x+12 5/x-12 x-12/5

OpenStudy (anonymous):

By f-1(x) you mean the inverse I assume. To do that just swap the values of your function with y<->x. y=5x+12 becomes x=5y+12 and now solve for y.

OpenStudy (anonymous):

wut? this is confusing

OpenStudy (anonymous):

Do you know how to solve the equation for y, that means to rearrange the equation I have given, x=5y+12 to have y alone and all the numbers on x's side.

OpenStudy (anonymous):

let me see,

OpenStudy (anonymous):

one moment

OpenStudy (anonymous):

y=5x+12?

OpenStudy (anonymous):

isnt that what is up their ??

OpenStudy (anonymous):

No man.. you need to do algebraic operations. To get 12 on the other side, you need to subtract 12 from both sides, so you will have, x-12 =5y Now what do you do with the 5 that is multiplying y. How do you move that from the right side to the left side?

OpenStudy (anonymous):

oh man this is tuff, i remember learning this last year.

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

Dude, this is as easy as math gets.. It aint tuff.. trust me.

OpenStudy (anonymous):

i give up, though, I was thinking this, 5y+x=12

OpenStudy (anonymous):

What grade are you on?

OpenStudy (anonymous):

junior, but I failed segment 2 of algebra in my sophomore year bc i slacked off

OpenStudy (anonymous):

i guess you have the right to tell me im dumb

OpenStudy (anonymous):

no dude you're not dumb, I just wanted to know how to explain the problem to you. Alright, check this out. x-12=5y You divide both sides by 5 because that number is multiplying y on the right side. So when you divide both sides you will cancel out the one that's multiplying y and. Therefore your equation now becomes, x/5 - 12/5 = y That is the inverse of your initial function

OpenStudy (anonymous):

oh, i get it

OpenStudy (anonymous):

how would i solve for this question?

OpenStudy (anonymous):

Are those the only possibilities that you have?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

None of them are correct.

OpenStudy (anonymous):

which one would be the closest?

OpenStudy (anonymous):

if you had something like 1/5x - 12/5, that would be the closest

OpenStudy (anonymous):

ill just guess. Thanks for your help though!

OpenStudy (anonymous):

lol, sure you're welcome.